OpenStudy (anonymous):

ack took 5.5 seconds to fall down the hill. His initial velocity was 0.5 meters per second. His final velocity was 2.5 meters per second. Calculate his average acceleration. I chose the answer 2.0 m/s/s is that correct?

OpenStudy (ajprincess):

use v=u+at 2.5=0.5+a*5.5 2.5-0.5=5.5a 2/5.5=a

OpenStudy (ajprincess):

is it a multiple choice question?

OpenStudy (ajprincess):

@13cscazalar

OpenStudy (anonymous):

no its not

OpenStudy (ajprincess):

Then hw did u choose ur answer 2m/s/s?

OpenStudy (anonymous):

Thanks the equation worked

OpenStudy (ajprincess):

yw:)